ddawgg:

A football player kicks a 0.94 kg football with a force of 2.4 N. Calculate the acceleration of the football as the player kicks the ball. Round your answer to the nearest tenth of a m/s2.

jimthompson5910:

We use Newton's second law F = m*a In this case, F = 2.4 newtons = force applied m = 0.94 kg = mass a = unknown acceleration

jimthompson5910:

So you need to solve for 'a' F = m*a 2.4 = 0.94*a

ddawgg:

whats the answer\

jimthompson5910:

Divide both sides by 0.94 to solve for 'a'

ddawgg:

thx merry christmas

jimthompson5910:

No problem

ddawgg:

so divide 2.4 by 0.94

jimthompson5910:

correct

ddawgg:

2.5531914

ddawgg:

rounded is 2.6

jimthompson5910:

You have the correct answer

ddawgg:

thank you very much

jimthompson5910:

You're welcome

ddawgg:

here is another one A model rocket accelerates at 15.3 m/s2 with a force of 44 N. Calculate the mass of the rocket. Round your answer to the nearest tenth of a kg. The mass of the rocket is kg.

ddawgg:

the massof the rocket is _______kg.

jimthompson5910:

So we have a = 15.3 and F = 44

jimthompson5910:

Solve for m F = m*a 44 = m*15.3 To solve for m, we need to divide both sides by 15.3 to undo the multiplication.

ddawgg:

2.9

jimthompson5910:

Correct answer

ddawgg:

thx bud

jimthompson5910:

Sure thing

ddawgg:

edginuity sucks

ddawgg:

What is the mass of a crate if a force of 200 N causes it to accelerate at 8 m/s2? (Formula: F=ma)

jimthompson5910:

So we have F = 200 and a = 8 this time

ddawgg:

i divid 200 by 8

jimthompson5910:

Yes those are the correct steps. What does that lead to?

ddawgg:

25 kg

jimthompson5910:

correct

ddawgg:

you are really good

jimthompson5910:

Thanks I appreciate it

ddawgg:

no thanks you

jimthompson5910:

sure thing

ddawgg:

Luis and Aisha conducted an experiment. They exerted different forces on four objects. Their results are shown in the table. A 3-column table with 4 rows. The first column labeled Object has entries 1, 2, 3, 4. The second column labeled Mass has entries 10 kilograms, 100 grams, 10 grams, 1 kilogram. The third column labeled Force (Newtons) has entries 4 Newtons, 20 Newtons, 4 Newtons, 20 Newtons. Based on the data, which object has the greatest acceleration?

ddawgg:

i described the table

ddawgg:

well it described the table

jimthompson5910:

it would be better to post a screenshot

ddawgg:

how o you do that on a chroam book

jimthompson5910:

I'm not sure I never used them before. This site might help https://www.digitaltrends.com/computing/how-to-take-a-screenshot-on-a-chromebook/

ddawgg:

1 attachment
jimthompson5910:

Thanks

ddawgg:

np

jimthompson5910:

So what you'll do is divide each force over its corresponding mass. Then you'll pick the largest result.

jimthompson5910:

For row 3, you need to convert 10 grams to kilograms

jimthompson5910:

1 kg = 1000 grams 1 gram = 0.001 kg 10 grams = 0.01 kg

jimthompson5910:

Row 2 is a similar story, 100 grams = 0.1 kg

ddawgg:

A ball has a mass of 140 g. What is the force needed to accelerate the ball at 25 m/s2? (Formula: F=ma)

jimthompson5910:

F = ma F = 0.140*25 F = ?? Note how I converted 140 grams to 0.140 kg

ddawgg:

so divide 0.140 by 25

jimthompson5910:

No you compute 0.140*25 the * means multiply

ddawgg:

3.500000000000...

jimthompson5910:

3.5 yes

ddawgg:

1 attachment
jimthompson5910:

Your thoughts?

jimthompson5910:

hint: think of a tug of war

ddawgg:

1 attachment
ddawgg:

wrong one

ddawgg:

mb

jimthompson5910:

Not sure what you mean

ddawgg:

1 attachment
jimthompson5910:

Choice C is correct

ddawgg:

Laura sees a horse pulling a buggy. She wonders how it can accelerate if the action of the horse pulling the cart would cause an equal and opposite reaction of the cart pulling on the horse. Which explanation best answers her question?

ddawgg:

a.The net forces exerted on the horse and cart are the same, but the action force is larger. b.The net forces exerted on the horse and cart are the same, but they are in the same direction. c.The net forces exerted on the horse and cart are not the same, so the two forces cancel out. d.The net forces exerted on the horse and cart are not the same, so they are not balanced forces.

jimthompson5910:

I'm not sure about this one

ddawgg:

ok how about

ddawgg:

1 attachment
jimthompson5910:

The smooth ice reduces friction, which in this case is the reactionary force. This is because it reacts to the force of the puck being pushed. For example, if the puck is pushed to the right, then the reactionary force friction pushes to the left to slow down the puck. 

jimthompson5910:



ddawgg:

so it slowes down the puck

jimthompson5910:

yes

jimthompson5910:

Smooth ice doesn't eliminate friction entirely

ddawgg:

ok then

jimthompson5910:

Friction is a good thing or else the players wouldn't be able to stop, slow down, or change direction.

ddawgg:

1 attachment
jimthompson5910:

Your thoughts?

jimthompson5910:

Hint: think of something that isn't moving and what it takes to get it moving

ddawgg:

i got it